OpenStudy (ledah):

graph of a line passing through the ordered pair negative four comma negative four and the ordered pair six comma one

OpenStudy (anonymous):

Your keyboard has keys on it specifically for typing numbers and commas, just so you know. (-4,-4) and (6,1)

OpenStudy (anonymous):

There are a number of ways you can do this. I like slope intercept form, y=mx+b where m is the slope and b is the y-intercept. Find the slope first. Rise over run.

OpenStudy (ledah):

I know about the comma's i just copied and pasted this :L

OpenStudy (anonymous):

Once you know m, pick a point you know (x,y), plug x, y, and m into y = mx + b, then solve for b to get the intercept. Use m and b to write the general form.

OpenStudy (ledah):

but please continue

OpenStudy (anonymous):

Haha, gotcha. I'm just messing with you =) Try that out and ask me any questions that come up.

OpenStudy (ledah):

y = ½x - 2 is this it?

OpenStudy (anonymous):

Great job =)

OpenStudy (ledah):

thanks

OpenStudy (anonymous):

\[m = \frac{\text{rise}}{\text{run}} = \frac{1-(-4)}{6-(-4)} = \frac{5}{10} = \frac{1}{2}\] Now to find the value of b, I'll use the point (6,1) \[y = mx + b \rightarrow 1 = \frac{1}{2}*6 + b \rightarrow 1 = 3 + b\\ \rightarrow 1-3 = -2 =b\] So m = 1/2 and b = -2 gives y = (1/2)x -2